OpenStudy (anonymous):

Solve 3x-5=10 and does the x=

OpenStudy (anonymous):

x=5

OpenStudy (anonymous):

add 5 on both sides you will get 3X = 10 +5 = 15 therefore x = 15/3 = 5

OpenStudy (anonymous):

good job suvesh253 thats right